OpenStudy (anonymous):

A triangle has side lengths of 7, 10, and 12. Is the triangle a right triangle? A) Yes B) No

Parth (parthkohli):

Apply the Pythagorean theorem here. Check if: \( \color{Black}{\Rightarrow 7^2 + 10^2 = 12^2}\)

OpenStudy (anonymous):

Then that's no, because it comes out to be 293.. A right triangle is nly 90

OpenStudy (anonymous):

only*

Parth (parthkohli):

No, we're talking about the side lengths here, not the angle.

OpenStudy (anonymous):

Huh wha is the side length of a triangle?

Parth (parthkohli):

In a right triangle: \( \color{Black}{\Rightarrow a^2 + b^2 = c^2 }\) c is the longest side aka hypotenuse.

OpenStudy (anonymous):

Ok.. I still don't undestand the question.

Parth (parthkohli):

You have to check if \(7^2 + 10^2 = 12^2\) for checking if it is a right triangle.

OpenStudy (anonymous):

I did, 7²=49 , 10²=100, 13²=144 && adding them all together you get 293

Parth (parthkohli):

No! Why are you adding?

OpenStudy (anonymous):

Why are you getting mad?

Parth (parthkohli):

\(\Large \textbf{LHS}\) \( \color{Black}{\Rightarrow 7^2 + 10^2 }\) \( \color{Black}{\Rightarrow 49 + 100 }\) \( \color{Black}{\Rightarrow 149 }\) \(\Large \textbf{RHS}\) \( \color{Black}{\Rightarrow 12^2 }\) \( \color{Black}{\Rightarrow144 }\) Is 149 equal to 144?

OpenStudy (anonymous):

No it's not equal..

OpenStudy (anonymous):

149 is greater

Parth (parthkohli):

Then these are not the sides of a right angled triangle.

OpenStudy (anonymous):

k. thanks